When a solid dissolves in water, the solution may become hotter or colder. The dissolution enthalpy (dissolving) can be determined using a coffee cup calorimeter. In the laboratorv a general chemistry student finds that when 1.89 g CaBr2(s) is dissolved in 101.10 g water, the temperature of the solution increases from 24.47 to 26.81 °C Thermometer Cardboard or Styrofoam lid The heat capacitv of the calorimeter (sometimes referred to as the calorimeter constant) was determined in a separate experiment to be 1.55 J/oC Nested Styrofoam cups Reaction occurs in solution. Based on the students observation, calculate the dissolution enthalpy of CaBr2(s) in kJ/mol Assume the specific heat capacity of the solution is equal to the specific heat capacity of water. Алі, Н kJ/mol
Answer #1

|Q| = Ccal*delta T + m(water)*C(water)*delta T

= 1.55*(26.81-24.47) + 101.10*4.184*(26.81-24.47)

= 3.627 + 989.8

= 993.4 J

This is heat released by dissociation

So,

Q = -993.4 J

Molar mass of CaBr2,

MM = 1*MM(Ca) + 2*MM(Br)

= 1*40.08 + 2*79.9

= 199.88 g/mol

mass(CaBr2)= 1.89 g

use:

number of mol of CaBr2,

n = mass of CaBr2/molar mass of CaBr2

=(1.89 g)/(1.999*10^2 g/mol)

= 9.456*10^-3 mol

Now use:

delta H dis = Q/number of mol

= -993.4 J / 9.456*10^-3 mol

= -1.05*10^5 J/mol

= -105 KJ/mol

Answer: -105 KJ/mol